OpenStudy (anonymous):

Could someone help me solve: x/x+5 - 4 = 5/x+5

OpenStudy (anonymous):

please put () in your question so I know how to answer it ((x/x)+5 or x/(x+5), for example?)

OpenStudy (anonymous):

sorry, \[\frac{ x }{ x+5 } - 4 = \frac{ 5 }{ x+5 }\]

OpenStudy (anonymous):

multiply -4 by (x+5)/(x+5) makes it -(4x+20)/(x+5), remove all x+5 from denominators, and you have x-4x-20=5

OpenStudy (anonymous):

then solve from there

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

I don't understand why one would just remove all x+5 from the denominators, could you explain it?

OpenStudy (anonymous):

if the denominator of every term in an equation is equal, one can simply cross them all out. this is the same as multiplying each term in the equation by the denominator- it effectively causes them to cease to exist and save you the trouble of having tons of fractions in your equation.

OpenStudy (anonymous):

so you're really multiplying both sides of the equation by x+5, which gets rid of the denominators.

OpenStudy (anonymous):

ok, I get it, thanks :)

OpenStudy (anonymous):

np